Circular No. 6482

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

X0115+634 AND CYGNUS X-3 R. Remillard, Massachusetts Institute of Technology (MIT); and B.

Thomas and T. Takeshima, Goddard Space Flight Center (GSFC), report for the RXTE ASM Team at MIT and GSFC: “The Be binary/pulsar X0115+634 is again increasing in x-ray brightness. Beginning on Sept. 26, the RXTE All Sky Monitor began detecting the source, with an average flux of 15 ± 3 mCrab at 2–12 keV on that day. By Sept. 30, the average flux had increased to 40 ± 3 mCrab. Cygnus X-3 has also displayed increasing brightness. During Mar.–Aug., the 2- to 12-keV flux usually varied between 70 and 140 mCrab. However, during Sept. most of the average daily measurements were between 140 and 230 mCrab, with increasing variability and spectral softening. On Sept. 28 and 29, the average flux was 275 ± 5 mCrab — while on Sept. 30, the flux dropped to 146 ± 3 mCrab with continued variability on timescales of hours.”

V694 MONOCEROTIS = MWC 560 L. Georgiev, R. Barba, and G. Koenigsberger, Instituto de Astrono-

mia, Universidad Nacional Autonoma de Mexico; and T. Tomov, Institute of Astronomy, Bulgarian Academy of Sciences, write: “Echelle spectra of the peculiar stellar object V694 Mon = MWC 560 (α = 7^{h}25^{m}51^{s}, δ = −7^{o}44′.1, equinox 2000.0), obtained with the 2.1-m telescope of the National Observatory in San Pedro Martir, during Sept. 16-19 show similar features to those observed during the last reported active periods of MWC 560 in 1990 Jan. (IAUC 4955) and 1991 Oct. (IAUC 5354, 5355). Highly blueshifted and variable absorption is present in H I, He I, Na I, and Ca II, but only in multiplet 42 of Fe II. Blue-edge velocities for Hβ are as follows: Sept. 16.5 UT, −2030 km/s; 17.5, −2200; 18.5, −2540; 19.5, −3400. A new active phase appears to be starting.”

COMET C/1995 O1 (HALE-BOPP) Visual m_{1} and coma-diameter estimates (B = binoculars): Sept. 2.86

UT, 5.8, 13′ (B. H. Granslo, Fjellhamar, Norway, 10×50 B); 9.14, 5.5, – (R. Keen, Mt. Thorodin, CO, naked eye); 13.11, 5.7, 13′ (G. W. Kronk, Troy, IL, 7×35 B); 13.83, 5.4, – (M. V. Zanotta, St. Barthelemy, Italy, naked eye); 16.11, 5.6, 12′ (A. Hale, Cloudcroft, NM, 10×50 B); 21.83, 5.2, 17′ (A. Pereira, Carnaxide, Portugal, 9×34 B); 30.10, 5.5, 13′ (Hale); Oct. 1.84, 5.3, 14′ (C. Vitorino, Cabo da Roca, Portugal, 9×34 B).

1996 October 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green